The sum of the digits of a two-digit number is 14. The tens digit is the square of a number which is 2 less than the units digit. Find the two-digit number.
Found 3 solutions by mananth, greenestamps, ikleyn:
Answer by mananth(16946)

The sum of the digits of a two-digit number is 14. The tens digit is the square of a number which is 2 less than the units digit. Find the two-digit number.
Let the digits in tens place be x
and the digit in units place be y
x+y=14............................................(1)
The tens digit is the square of a number which is 2 less than the units digit.
y= (x-2)^2.........................................(2)
substitute y in (1)
x+(x-2)^2= 14
x+x^2-4x+4=14
x^2-3x-10=0
factorize
x^2-5x+2x-10=0
x(x-5)+2(x-5)=0
(x-2)(x-5)=0
x=2 or x=5
x cannot be equal to 2
so x=5
y=9
The number = 59



Answer by greenestamps(13198)



In the response from the other tutor, she mixed up her variables, ending with an answer that doesn't satisfy the given conditions.

Her process is fine; you can turn her solution into a valid and correct solution by keeping the variables straight.

While undoubtedly a formal algebraic solution was wanted, the problem is solved quickly and easily using logical reasoning:

The sum of the two digits is 14, so the tens digit can only be 5, 6, 7, 8, or 9.

Since the tens digit is a perfect square, it must be 9.

That makes the units digit 14-9 = 5.

ANSWER: 95


Answer by ikleyn(52776)

.

In this problem, the sum of two digits 14 can be produced by the only one way: 14 = 9 (as the ten's digit, which is a square) + 5.

So, the all other info is simply not needed.